OpenStudy (anonymous):

What number would you put in a matrix to represent a term that is missing in the original equation? A. 0 B. 1 C. 2 D. -1

OpenStudy (dan815):

ah look at this

OpenStudy (dan815):

say you have the equations 3x + 2y + z = 0 -y + 2z = 2 x +2y = 1

OpenStudy (dan815):

how can you write this as a matrix

OpenStudy (dan815):

have no x is the same thing as saying 0 * x

OpenStudy (dan815):

3x + 2y + z = 0 0x -y + 2z = 2 x +2y + 0z = 1

OpenStudy (anonymous):

well i was think ... if you have no x or y or a number nothing is there right ? so it'll be 0

OpenStudy (dan815):